Calculate the mean for the population data. Select one random sample of size three and calculate the sample mean Compute the sampling error?


1
Expert's answer
2021-12-27T14:08:45-0500

Solution:

Data is missing. We assume a population consists of three numbers 2, 5, 8.

Population mean = (2+5+8)/3=15/3 = 5.

All possible samples of size 2:{(2,5), (2,8), (5,8)}.

Sample means for these samples: 3.5, 5, 6.5.

Mean of the sample means: (3.5+5+6.5)/3=5.

So, the sample mean is an unbiased estimate of the population mean.

Sampling error=Population mean - sample mean

= 5-5

= 0